# Equipping Sketch Patches with Context-Aware Positional Encoding for Graphic Sketch Representation

### v1

@@ Abstract
The drawing order of a sketch records how it is created stroke-by-stroke by a human being. For graphic sketch representation learning, recent studies have injected sketch drawing orders into graph edge construction by linking each patch to another in accordance to a temporal-based nearest neighboring strategy. However, such constructed graph edges may be unreliable, since a sketch could have variants of drawings. In this paper, we propose a variant-drawing-protected method by equipping sketch patches with context-aware positional encoding (PE) to make better use of drawing orders for learning graphic sketch representation. Instead of injecting sketch drawings into graph edges, we embed these sequential information into graph nodes only. More specifically, each patch embedding is equipped with a sinusoidal absolute PE to highlight the sequential position in the drawing order. And its neighboring patches, ranked by the values of self-attention scores between patch embeddings, are equipped with learnable relative PEs to restore the contextual positions within a neighborhood. During message aggregation via graph convolutional networks, a node receives both semantic contents from patch embeddings and contextual patterns from PEs by its neighbors, arriving at drawing-order-enhanced sketch representations. Experimental results indicate that our method significantly improves sketch healing and controllable sketch synthesis.
@@
Graphic sketch representation , positional encoding , graph convolutional network , controllable sketch synthesis , sketch healing
@@ 1 Introduction
Free-hand sketches carry vivid emotions and messages for communication through human history. The drawing order of a sketch records how it is drawn stroke-by-stroke by a human being. It provides models a unique view to review how sketch components are organized by a sequential positional structure, which cannot be stored on the canvas by sketch images seen in daily life.
@@
Figure 1: Two approaches to inject drawing orders into graphic sketch representation, when dealing with variants of sketch drawings. (a) Constructing graph edges by drawing orders as in [1, 2], i.e., linking the neighboring sketch components by a drawing order. (b) The proposed approach. We inject drawing orders into PEs, which only enrich graph nodes but never attend the edge construction.
@@
Recent studies have utilized the sequential information from sketch drawing orders to cooperate with the visual patterns from sketch images for representation learning. One common approach is employing a two-branch encoder to capture a pair of latent codes from a sketch image and its corresponding sketch sequence simultaneously, and further combine them for the final sketch representation [3, 4, 5, 6]. Another way is injecting drawing orders into graphic sketch representation by linking sketch patches according to drawing-order-based temporal proximity [1, 2]. The hybrid of temporal and visual features via sketch graphs is flexible and beneficial for learning accurate sketch representation.
Due to variations of sketch drawings, it is doubtful whether the sequential information stored in drawing orders could reveal reliable relationships between sketch components, e.g., constructing graph edges as in [1, 2]. For example, as shown in Fig. 1, when drawing a pig face, its nose and ears could be drawn next to each other in order A, but are separated in B, which arrives at two different groups of graph edges with various sequential relationships among sketch components, shown in Fig. 1(a). If graph edges are constructed by such variable relationships, the passing messages over them for message aggregation via graph convolutional network (GCN) layers [7] might dilute or interrupt the target node. As a result, though no additional information from drawing order attends the representation learning, SP-gra2seq [8], which constructs graph edges in a more reliable way by semantic similarity, still outperforms SketchHealer [1] and SketchHealer 2.0 [2] on sketch representation learning.
@@
It motivates us to make better use of sketch drawing orders by injecting them into graph nodes, shown in Fig. 1(b). More specifically, each sketch patch regarded as graph nodes are enriched by its sequential position encoded by a vectorized positional encoding (PE) [9, 10]. PEs record when a patch is drawn in a drawing order [^1] The cropping centers of sketch patches are selected following the drawing orders [1, 2, 8]. Thus, we transplant the sequential positions in drawing order to rank patches.. We borrow the term context from the field of linguistics to describe the neighboring patches which surround a focal patch in a drawing order. By the attendance of PEs, the position-unaware graph nodes are equipped with contextual information. During the message aggregation via GCN layers, each node receives not only the content of neighboring nodes but also when they are drawn in the timeline, guiding our method to capture more hidden patterns from the unique view of drawing. Moreover, our graph edges are constructed by semantic similarity between patches as in [8], leaving sequential positions away from node linking to protect graph construction from variants of sketch drawings.
@@
To realize the above proposals, we propose Drawing-order-enhanced Context-aware graph to sequence (DC-gra2seq) to equip graphic sketch representation with drawing orders. Each patch embedding, captured by a convolutional neural network (CNN) encoder, is equipped with a sinusoidal absolute PE to highlight the sequential position in drawing orders. Moreover, its neighboring patches, which are ranked by the values of self-attention scores between patch embeddings, are offered with learnable relative PEs to restore the contextual positions within a neighborhood. Both patch embeddings and PEs are incorporated by a GCN encoder to obtain the final sketch representation, which is sent into a recurrent neural network (RNN) decoder for sketch generation. To summarize, we make the following contributions:
@@
    We propose DC-gra2seq to learn graphic sketch representation by introducing context-aware PE to make better use of sketch drawing orders.
@@
    We rethink the attendance of drawing orders in graphic sketch representation, and equip them on graph nodes to access the sequential positions from drawing orders, but keep them away from graph edges to protect sketch learning from variants of sketch drawings.
@@
Figure 2: Learning graphic representation $\bm{y}_{t}$ of sketch $\bm{S}_{t}$ by DC-gra2seq. The cropped sketch patches $\{\bm{p}_{tm}\}$ along with the resized full sketch $\bm{p}_{t0}$ are embedded by a CNN encoder as patch embeddings $\{\bm{v}_{tm}\}$. The absolute PE $\bm{P}$ embedded by original sketch drawing order and the relative PE $\bm{R}$ restoring contextual positions between patches are incorporated with $\{\bm{v}_{tm}\}$, weighted by the masked self-attention scores $\bm{A}_{t}$ computed by solo $\{\bm{v}_{tm}\}$. A GCN layer collects the aggregated messages to produce the final sketch code $\bm{y}_{t}$.
@@ 2.1 Learning sketch representations
Sketches are usually formed by raster images to restore visual patterns from the canvas, or formed by sequences of coordinates to record how they are drawn stroke-by-stroke by human. For sketch images, CNN-based methods, e.g., [11, 12, 13, 14], could be utilized for representation learning to capture the spatial dependencies among drawing strokes. For sketch sequences, in order to extract temporal relationships between strokes, RNN-based, e.g., [15, 16], or transformer-based methods, e.g., [17, 18] have been proposed to learn the contextual information along sketch drawing orders. Moreover, using dual-branch encoders, e.g., [3, 4, 6, 5], is beneficial to learn accurate and efficient latent representations from paired sketches (a sketch image and its corresponding sketch sequence), by making better use of both temporal and spatial relationships among sketch components.
Recently, graphic sketch representation is demonstrated as effective ways for sketch representation learning in sketch synthesis [1, 2, 19, 8], sketch recognition [20, 21, 22], sketch segmentation [23, 24] and sketch-based image retrieval [25]. A sketch is formed as a graph with multiple components, whose captured features are aggregated by graph neural network (GNN) [26] or GCN layers to obtain the final sketch representation. Graph nodes can be sketch patches cropped from sketch images [1, 2, 8], coordinates on the latticed canvas [19], etc, which carry local sketch patterns. Graph edges can be constructed by linking sketch components to follow sketch drawing order [1, 2], Euclidean distances between sketch components [19, 24], or semantic similarity [8], etc. Graphic representation enables the message passing between sketch components for learning better sketch representation.
We aim to equip graph nodes with contextual positions from sketch drawing order via PEs, which makes message aggregating process via GCN layers context-aware for learning accurate sketch representations.
@@ 2.2 Positional encoding
Positional encoding makes better use of some valuable sequential or structural information, when training position-unaware networks, e.g., self-attention network [10] and graph attention network (GAT) [27]. PEs give these architectures a sense of which portion of the sequence in input (or output) it is currently dealing with. Recent studies embedded the positional information by absolute [9, 10], relative PEs [28, 29, 30] or both [31]. Absolute PEs restore the absolute positions of the input tokens, indicating where a specific token is located in a sequence or a canvas. Relative PEs encode the relative pairwise distances or positions of a pair of tokens, which is invariant to the position of focal token in contrast to absolute PE. The positional information can be embedded by either fixed [10, 31] or learnable PEs [9, 28, 30, 32, 33].
We inject sketch drawing orders into graph nodes in graphic sketch representation by equipping sketch patches with both absolute and relative PEs. Each patch collects the information from its neighboring patches along with their corresponding PEs during a message aggregating process.
@@ 3 Methodology
Fig. 2 offers an overview to indicate how DC-gra2seq learns graphic sketch representation. The embeddings of the cropped sketch patches as graph nodes are captured by a CNN encoder, and the graph edges between them are constructed by the masked self-attention scores computed from solo patch embeddings. During message aggregation via a GCN layer, both absolute and relative PEs are equipped on graph node to make use of the sequential and contextual positions in drawing orders. And the final sketch code is fed into an RNN decoder to reconstruct the input sketch in a sequence form.
3.1 Injecting sketch drawing order into PE
For the $t$-th sketch image $\bm{S}_{t}\in\mathbb{R}^{640\times 640}$ as input, we crop $M$ sketch patches $\{\bm{p}_{tm}\}_{m=1}^{M}$ with the size of $256\times 256$ from the canvas. We adopt the selecting rule utilized in [1, 2, 8] to determine the cropping centers of patches, which arranges these patches in a sequential order in accordance to the drawing orders. More specifically, as the cropping center of $\bm{p}_{tm}$ is located on the stroke drawn before the one of $\bm{p}_{t,m+1}$, $\bm{p}_{tm}$ is positioned prior to $\bm{p}_{t,m+1}$ in the sequential order. Such sequential information makes patches position-aware, and we name the patches surrounding a focal one in the drawing order contexts of one another.
In order to make use of the sequential information, we firstly inject sketch drawing orders into sinusoidal absolute PE [10], shown in Eq. (3.1).
@@ 3.1 Injecting sketch drawing order into PE
Secondly, in order to encode the contextual distances between sketch patches along the drawing order, we introduce the learnable relative PE, $\bm{R}(i,j)$ to represent the contextual distance between patches $\bm{p}_{ti}$ and $\bm{p}_{tj}$. Each $\bm{R}(i,j)$ is a learnable vector with the same length of $\bm{P}(i)$, and we restrict $\bm{R}(i,j)$ with the following three conditions:
@@
$\bm{R}(i,j)$ $=\bm{R}(j,i),$ (3)
$\bm{R}(i,i)$ $=\bm{0}.$ (4)
Our relative PE is target-invariant in Eq. (2), i.e., a relative PE is only sensitive to the contextual distance between patches, but is invariant with the contents of patches. Two pairs of patches with the same contextual distance, e.g., ($\bm{p}_{ti}$, $\bm{p}_{t,i+k}$) and ($\bm{p}_{tj}$ and $\bm{p}_{t,j+k}$) share the same value of relative PE. Our relative PE is undirected in Eq. (3), i.e., it does not have orientations. Finally, we also fix the relative PE between one patch and itself at the token $\bm{0}$ with all elements valued by zero.
As a result, for a sketch represented by $M$ sketch patches, totally $M$ vectorized relative PEs are required in DC-gra2seq.
@@ 3.2 Context-aware graphic sketch representation
We link the sketch patches $\bm{p}_{ti}$ and $\bm{p}_{tj}$ to construct a graph edge by considering their self-attention score $\alpha^{\text{atten}}_{t}(i,j)$, computed by their patch embeddings $\bm{v}_{ti}$ and $\bm{v}_{tj}$. In practice, we implement the attention operation by using the cosine similarity, which is named as synonymous proximity in [8].
$\alpha^{\text{atten}}_{t}(i,j)=\frac{\bm{v}_{ti}^{\top}\bm{v}_{tj}}{\|\bm{v}_{ti}\|_{2}\cdot\|\bm{v}_{tj}\|_{2}},$ (5)
where $\|\cdot\|_{2}$ denotes the L2 norm. A large value of $\alpha^{\text{atten}}_{t}(i,j)$ indicates that the patch $\bm{p}_{ti}$ is highly relevant to $\bm{p}_{tj}$ in content. It is worth noting that neither the absolute PE nor relative PE attends the calculation of attention scores. The absence of PEs reduces the uncertainty brought by variants of sketch drawings, driving to reliable edge constructions for the passing message aggregation.
We utilize the computed self-attention score $\alpha^{\text{atten}}_{t}(i,j)$ in Eq. (5) to construct graph edges, and each node is only linked to the ones with the top-2 largest values among $\{\bm{v}_{tm}\}_{m\neq i}$. We arrive at the masked attention scores, stored in an adjacency matrix $\bm{A}_{t}$, with the element $\bm{A}_{t}(i,j)$ in $i$-th row and $j$-th column computed by,
$\bm{A}_{t}(i,j)=\left\{\begin{array}[]{ll}1,&j=i,\\ 0.5\cdot\alpha^{\text{atten}}_{t}(i,j),&j=j^{*},\\ &j^{*}=\mathop{\arg\max}\limits_{m\neq i}~{}\alpha^{\text{atten}}_{t}(i,m),\\ 0.2\cdot\alpha^{\text{atten}}_{t}(i,j),&j=j^{\prime},\\ &j^{\prime}=\mathop{\arg\max}\limits_{m\neq i,j^{*}}~{}\alpha^{\text{atten}}_{t}(i,m),\\ 0,&\text{otherwise.}\end{array}\right.$ (6)
@@
Eq. (7) presents the message aggregating process for node $\bm{v}_{ti}$ via a single GCN layer.
@@
where $\mathcal{N}(\bm{v}_{ti})$ denotes the neighborhood of $\bm{v}_{ti}$ determined by the masked self-attention scores in Eq. (6).
The target node $\bm{v}_{ti}$ receives the passing messages containing both the patch embedding $\bm{v}_{tj}$ and its corresponding relative PE $\bm{R}(i,j)$, from the neighbors. A neighboring node not only delivers its semantic content captured from sketch patch but also the contextual position between them. After incorporating all the passing message weighted by their attention score $\bm{A}_{t}(i,j)$, DC-gra2seq is able to learn what the neighboring nodes contain to benefit the target $\bm{v}_{ti}$ and to check how far their contextual distances are toward $\bm{v}_{ti}$. Moreover, we further incorporate $\bm{v}_{ti}$ with its absolute PE $\bm{P}(i)$ to highlight where $\bm{v}_{ti}$ is positioned in the drawing order. Especially, both absolute and relative PEs only participate in node aggregation to enrich the patch embedding with sequential and contextual positions, but are kept away from the edge construction in $\bm{A}_{t}$, driving to variant-drawing-protected sketch representation.

### v2

@@ Abstract
When benefiting graphic sketch representation with sketch drawing orders, recent studies have linked sketch patches as graph edges by drawing orders in accordance to a temporal-based nearest neighboring strategy. However, such constructed graph edges may be unreliable, since the contextual relationships between patches may be inconsistent with the sequential positions in drawing orders, due to variants of sketch drawings. In this paper, we propose a variant-drawing-protected method by equipping sketch patches with context-aware positional encoding (PE) to make better use of drawing orders for sketch learning. We introduce a sinusoidal absolute PE to embed the sequential positions in drawing orders, and a learnable relative PE to encode the unseen contextual relationships between patches. Both types of PEs never attend the construction of graph edges, but are injected into graph nodes to cooperate with the visual patterns captured from patches. After linking nodes by semantic proximity, during message aggregation via graph convolutional networks, each node receives both semantic features from patches and contextual information from PEs from its neighbors, which equips local patch patterns with global contextual information, further obtaining drawing-order-enhanced sketch representations. Experimental results indicate that our method significantly improves sketch healing and controllable sketch synthesis. The source codes could be found at https://github.com/SCZang/DC-gra2seq.
@@
Graphic sketch representation , sketch drawing order , contextual relationships, positional encoding , graph convolutional network
@@ 1 Introduction
Free-hand sketches carry vivid emotions and messages for communication through human history. The drawing order of a sketch records how it is drawn stroke-by-stroke by a human being. It provides models a unique view to review how sketch components are positioned in sequential orders, which fails to be stored on the canvas by sketch images seen in daily life.
@@
Figure 1: Two approaches to inject drawing orders into graphic sketch representation, when dealing with variants of sketch drawings. (a) Constructing graph edges by drawing orders as in [1, 2], by linking the neighboring sketch components following a drawing order. (b) The proposed approach by injecting drawing orders into graph edges, keeping drawing orders away from graph edge construction.
@@
Recent studies have utilized sequential information from sketch drawing orders to cooperate with the visual patterns from sketch images for sketch learning. One common approach is employing a two-branch encoder to learn sketch representations from sketch images and their corresponding paired sketch sequence simultaneously [3, 4, 5, 6]. The combination of temporal and visual features is flexible and beneficial to learn accurate sketch representation.
Another effective way is injecting drawing orders into graphic sketch representation, e.g., constructing graph edges by drawing-order-based sequential positions [1, 2]. However, as pointed out by [7], due to variations of sketch drawings, the uncertainty of graph edge construction brought by different drawing orders, shown in Fig. 1(a), may interrupt message aggregation via graph convolutional network (GCN) layers [8], resulting in inaccurate sketch learning. Thus, it is doubtful whether the sequential information stored in drawing orders could reveal reliable contextual relationships between sketch components. Such sequential relationships in sketches are different from the contextual relationships between tokens in natural languages processing. Exchanging the positions of two tokenized phrases would corrupt a sentence to lose its semantic meaning. But for a sketch sequence, its drawing order never determines what it looks like on the canvas. As a result, the sequential positions in a drawing order might fail to precisely reveal their actual contextual relationships.
@@
It motivates us to make better use of sketch drawing orders by injecting them into graph nodes instead of edges, shown in Fig. 1(b). More specifically, each sketch patch regarded as graph nodes are enriched by its sequential position in drawing orders encoded by a vectorized absolute positional encoding (PE) [9, 10]. Absolute PE records when a patch is drawn in a drawing order [^1] The cropping centers of sketch patches are selected following the drawing orders [1, 2, 7]. Thus, we transplant the sequential positions in drawing order to rank patches.. Moreover, in order to capture some hints about the unseen contextual relationships, we embed the “contextual distances” between semantically related patches along drawing orders, by introducing a learnable relative PE. With the equipment of relative PE, each patch enables to realize how far away its semantically related patches are in a drawing order. By the attendance of PEs, during message aggregation via GCN layers, the position-unaware graph nodes receives not only the visual contents from patches but also their contextual information, guiding our method to capture more hidden patterns from the unique view of drawing. Furthermore, our graph edges are constructed by semantic similarity between patches as in [7], leaving drawing orders away from node linking to protect graph construction from variants of sketch drawings.
@@
To realize the above proposals, we propose Drawing-order-enhanced Context-aware graph to sequence (DC-gra2seq) to enhance graphic sketch representation learning with some contextual information from drawing orders. Each patch embedding, captured by a convolutional neural network (CNN) encoder, is equipped with a sinusoidal absolute PE to highlight the sequential position in drawing orders. Moreover, its neighboring patches, which are ranked by the values of cosine similarity between patch embeddings, are offered with learnable relative PEs to restore the contextual positions within a neighborhood. Both patch embeddings and PEs are incorporated by a GCN encoder to obtain the final sketch representation, which is sent into a recurrent neural network (RNN) decoder for sketch generation. To summarize, we make the following contributions:
@@
    We propose DC-gra2seq to learn graphic sketch representation by introducing context-aware PEs to make better use of sketch drawing orders.
@@
    We rethink the attendance of drawing orders in graphic sketch representation, and equip them on graph nodes to access some contextual information from drawing orders, but keep them away from graph edges to protect sketch learning from variants of sketch drawings.
@@ 2.1 Learning sketch representations
Sketches are usually formed by raster images to restore visual patterns from the canvas, or formed by sequences of coordinates to record how they are drawn stroke-by-stroke by human. For sketch images, CNN-based methods, e.g., [11, 12, 13, 14, 15], could be utilized for representation learning to capture the spatial dependencies among drawing strokes. For sketch sequences, in order to extract temporal relationships between strokes, RNN-based, e.g., [16, 17], or transformer-based methods, e.g., [18, 19] have been proposed to learn the contextual information along sketch drawing orders. Moreover, using dual-branch encoders, e.g., [3, 4, 6, 5], is beneficial to learn accurate and efficient latent representations from paired sketches (a sketch image and its corresponding sketch sequence), by making better use of both temporal and spatial relationships among sketch components.
Recently, graphic sketch representation is demonstrated as an effective way for sketch representation learning in sketch synthesis [1, 2, 20, 7], sketch recognition [21, 22, 23], sketch segmentation [24, 25] and sketch-based image retrieval [26]. A sketch is formed as a graph with multiple components regarded as graph nodes, whose captured features are aggregated by graph neural network (GNN) [27] or GCN layers to obtain the final sketch representation. Graph nodes can be patches cropped from sketch images [1, 2, 7], coordinates on a latticed canvas [20], etc, carrying local sketch patterns. Graph edges can be constructed by linking sketch components to follow sketch drawing order [1, 2], Euclidean distances between sketch components [20, 25], or semantic similarity [7], etc. Graphic representation enables the message passing between sketch components for learning better sketch representation.
It is beneficial to highlight some valuable information in graph edges or nodes to assist graphic representation learning, e.g., constructing graph edges by the spatial relationships between image regions in a criss-cross way [28]. We aim to equip graph nodes with contextual relationships from sketch drawing order via PEs, which makes message aggregating process via GCN layers context-aware for learning accurate sketch representations.
@@ 2.2 Positional encoding
Positional encoding makes better use of some valuable sequential or structural information, when training position-unaware networks, e.g., self-attention network [10] and graph attention network [29]. PEs give these architectures a sense of which portion of the sequence in input (or output) it is currently dealing with. Recent studies embedded the positional information by absolute [9, 10], relative PEs [30, 31, 32] or both [33]. Absolute PEs restore the absolute positions of the input tokens, indicating where a specific token is located in a sequence or a canvas. Relative PEs encode the relative pairwise distances or positions of a pair of tokens, which is invariant to the position of focal token in contrast to absolute PE. The positional information can be embedded by either fixed [10, 33] or learnable PEs [9, 30, 32].
We inject sketch drawing orders into graph nodes by equipping sketch patches with both absolute and relative PEs. Each patch collects the visual contents from its neighboring patches along with their contextual relationships in drawing orders, during a message aggregating process.
Figure 2: Learning graphic representation $\bm{y}_{t}$ of sketch $\bm{S}_{t}$ by DC-gra2seq. The cropped sketch patches $\{\bm{p}_{tm}\}$ along with the resized full sketch $\bm{p}_{t0}$ are embedded by a CNN encoder as patch embeddings $\{\bm{v}_{tm}\}$. The absolute PE $\bm{P}$ restoring sketch drawing order and the relative PE $\bm{R}$ encoding contextual relationships between patches are incorporated with $\{\bm{v}_{tm}\}$, weighted by masked coefficients $\bm{A}_{t}$ computed from $\{\bm{v}_{tm}\}$ only. A GCN layer collects all the information from $\bm{v}$, $\bm{P}$ and $\bm{R}$ to produce a final sketch code $\bm{y}_{t}$.
@@ 3 Methodology
Fig. 2 offers an overview to indicate how DC-gra2seq learns graphic sketch representation. The embeddings of cropped sketch patches as graph nodes are captured by a CNN encoder, and the graph edges between them are constructed by masked coefficients computed from patch embeddings. During message aggregation via a GCN layer, both absolute and relative PEs are equipped on graph node to make use of the sequential and contextual relationships in drawing orders, producing the final sketch code, which is fed into an RNN decoder to reconstruct the input sketch in a sequence form.
3.1 Encoding contextual information from sketch drawing order by PEs
For the $t$-th sketch image $\bm{S}_{t}\in\mathbb{R}^{640\times 640}$ as input, we crop $M$ sketch patches $\{\bm{p}_{tm}\}_{m=1}^{M}$ with the size of $256\times 256$ from the canvas. We adopt the selecting rule utilized in [1, 2, 7] to determine the cropping centers of patches, which arranges these patches in a sequential order in accordance to the drawing orders. More specifically, as the cropping center of $\bm{p}_{tm}$ is located on the stroke drawn before the one of $\bm{p}_{t,m+1}$, $\bm{p}_{tm}$ is positioned prior to $\bm{p}_{t,m+1}$ in the sequential order. Such sequential positions make patches position-aware, and we inject them into sinusoidal absolute PE [10], shown in Eq. (3.1).
@@ 3.1 Encoding contextual information from sketch drawing order by PEs
Furthermore, as mentioned in Sect. 1, the contextual relationships between drawing strokes may fail to be revealed by the sequential positions in drawing orders, due to variants of sketch drawings. Thus, we introduce the learnable relative PE $\bm{R}(i,j)$ to capture the contextual distances between sketch patches along drawing orders. $\bm{R}(i,j)$ is a learnable vector with the same length of $\bm{P}(i)$, encoding the contextual relationships between patches $\bm{p}_{ti}$ and $\bm{p}_{tj}$. We restrict $\bm{R}(i,j)$ with the following two conditions:
@@
$\bm{R}(i,j)$ $=\bm{R}(j,i).$ (3)
Firstly, our relative PE is target-invariant by Eq. (2), i.e., relative PE is insensitive to the content of patches. Two pairs of patches with the same contextual distances, e.g., ($\bm{p}_{ti}$, $\bm{p}_{t,i+k}$) and ($\bm{p}_{tj}$ and $\bm{p}_{t,j+k}$), share the same value of relative PE. Secondly, our relative PE is undirected by Eq. (3), i.e., the sequential order of two patches are not counted. $\bm{R}(i,j)$ only encodes how far away two patches $\bm{p}_{ti}$ and $\bm{p}_{tj}$ are in a drawing order. In addition, we make relative PE trainable to encourage DC-gra2seq to flexibly capture the generalized unseen contextual information among patches during training. And for a sketch represented by $M$ sketch patches, totally $M$ vectorized relative PEs are required in DC-gra2seq.
We simultaneously employ two types of PEs to juice both the obtainable sequential positions and the hidden contextual relationships between sketch patches from drawing orders. And PEs could be easily equipped on graph nodes and kept away from the construction of graph edges during DC-gra2seq training.
@@ 3.2 Context-aware graphic sketch representation
We link sketch patches $\bm{p}_{ti}$ and $\bm{p}_{tj}$ to construct a graph edge by considering their synonymous proximity [7], computed by the cosine similarity between their captured patch embeddings $\bm{v}_{ti}$ and $\bm{v}_{tj}$.
$\alpha_{t}(i,j)=\cos(\bm{v}_{ti},\bm{v}_{tj})=\frac{\bm{v}_{ti}^{\top}\bm{v}_{tj}}{\|\bm{v}_{ti}\|_{2}\cdot\|\bm{v}_{tj}\|_{2}},$ (4)
where $\|\cdot\|_{2}$ denotes the L2 norm. A large value of $\alpha_{t}(i,j)$ indicates that the patch $\bm{p}_{ti}$ is highly relevant to $\bm{p}_{tj}$ in content. It is worth noting that neither the absolute PE nor relative PE attends the calculation of $\alpha_{t}(i,j)$. The absence of PEs reduces the uncertainty in node linking brought by variants of sketch drawings, driving to reliable edge constructions for the passing message aggregation.
We utilize the computed $\alpha_{t}(i,j)$ in Eq. (4) to construct graph edges, and each node is only linked to the ones with the top-2 largest values among $\{\bm{v}_{tm}\}_{m\neq i}$. We arrive at the masked coefficients, stored in an adjacency matrix $\bm{A}_{t}$, with the element $\bm{A}_{t}(i,j)$ in $i$-th row and $j$-th column computed by,
$\bm{A}_{t}(i,j)=\left\{\begin{array}[]{ll}1,&j=i,\\ 0.5\cdot\alpha_{t}(i,j),&j=j^{*},\\ &j^{*}=\mathop{\arg\max}\limits_{m\neq i}~{}\alpha_{t}(i,m),\\ 0.2\cdot\alpha_{t}(i,j),&j=j^{\prime},\\ &j^{\prime}=\mathop{\arg\max}\limits_{m\neq i,j^{*}}~{}\alpha_{t}(i,m),\\ 0,&\text{otherwise.}\end{array}\right.$ (5)
@@
With the constructed graph edges stored in $\bm{A}_{t}$, we activate the message aggregating process for learning node $\bm{v}_{ti}$ via a GCN layer, shown in Eq. (6).
@@
where $\mathcal{N}(\bm{v}_{ti})$ denotes the neighborhood of $\bm{v}_{ti}$ determined by the adjacency matrix $\bm{A}_{t}$.
As the adjacency matrix $\bm{A}_{t}$ is sparse, a target node $\bm{v}_{ti}$ only receives the messages passing from two neighboring nodes, whose sketch patches are the most relevant in semantics. The neighboring node $\bm{v}_{tj}$ delivers not only its captured patch features $\bm{v}_{tj}$ but also the corresponding contextual relationships embedded by $\bm{R}(i,j)$. It encourages DC-gra2seq to realize what semantically related features a neighboring node has, and how far away the node is toward $\bm{v}_{ti}$ along the drawing order, equipping local patch patterns with contextual information observed from a global view. In addition, after incorporating all the passing messages, we provide each node with its sequential positions in the drawing order by $\bm{P}(i)$, to further enrich the visual patterns with temporal information from the drawing orders. It is worth noting that both absolute and relative PEs only participate in node aggregation, but are kept away from the edge construction in $\bm{A}_{t}$, driving to variant-drawing-protected sketch representation.